# Euclid Quick Data Release (Q1)

### v2

@@ 5.6 Comparing to other classifiers
Srsic and PSF Flux Measurements
@@
To benchmark our non-parametric inpainting method against a more traditional approach, we construct a classifier based on parametric modelling of galaxy light profiles. We utilise morphological parameters sourced directly from the Euclid MER Q1 catalogue (Euclid Collaboration: Romelli et al. 2025). Specifically, we use the provided flux_vis_sersic_euclid and flux_vis_psf_euclid measurements. These parameters are derived by the MER pipeline using the SourceExtractor++ software (Bertin et al. 2020; Kümmel et al. 2020). For each detected source, the pipeline performs two separate and independent model fits. The flux_vis_psf_euclid value represents the total flux obtained by fitting a model of the instrumental PSF to the source image. The flux_vis_sersic_euclid value represents the total flux from fitting a two-dimensional Srsic profile, which is designed to model the extended light distribution of a galaxy. The underlying assumption for a classifier based on these parameters is that sources with a dominant, unresolved central component (such as a QSO or a Type-1 AGN) will be well-described by a PSF model, resulting in a high flux_vis_psf_euclid value. Conversely, normal resolved galaxies will be better described by the extended Srsic profile, leading to a high flux_vis_sersic_euclid value. By comparing the relative strength of these two flux measurements (as shown in Fig. 15), we can construct a simple linear classifier to distinguish between point-like and extended sources. This baseline method is computationally inexpensive as it relies on pre-computed catalogue values, but it is fundamentally limited to this simple morphological distinction and may struggle with identifying AGN in complex host galaxies or at lower relative luminosities.
@@
The distribution of sources with these parameters is shown in Fig. 11. The equation of this linear boundary is y=2x1y=2x-1.
ROC Metric
@@
Our second comparison is with the original rate of change metric used in Sect. 4.3. As a clear distinction between the galaxy class and the others is visible, it is important to verify how such a metric performs as a classifier. This is especially true given that such a metric can be generated from the raw image, and no machine-learning model or precomputed values are required. The boundary for this classifier is $x=0.75$.
Performance against other candidates
@@
Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination. Therefore our analysis focuses primarily on achieving sufficiently high recall. For an initial candidate selection method, such as the one proposed here, the primary goal is often to compile a large, inclusive sample for potential follow-up observations or more detailed multi-wavelength analysis. In this context, high recall is typically prioritised over high precision, which measures the purity of the selected sample. The scientific cost of missing a potentially interesting source (a false negative) is often considered higher than the operational cost of including a contaminant (a false positive) that can be filtered out in subsequent analysis stages. Therefore, our primary goal is to demonstrate the capability of our model to achieve high recall across a wide variety of AGN and QSO candidates which have been selected via different, more conventional methods and physical mechanisms.
Precision and Recall are defined as follows
@@
For a given classification task (e.g., identifying DESI QSO objects):
@@
The test set scores for each respective selection are shown in Fig. 12, with the collated results for stars, galaxies, AGN and QSOs shown in Fig. 25. All predictions are made from images that were not used during training nor used in the creation of any of the decision boundaries for the classifiers. Due to differing constraints for each selection, we split the analysis to allow for a more representative comparison.
Galaxies
Our galaxy selection is determined by the DESI spectral type classification SPECTYPE=GALAXY (DESI Collaboration et al. 2024), with AGN contaminants removed following the procedure detailed in Euclid Collaboration: Matamoro Zatarain et al. (2025), hereafter called MZ25. Negative instances are formed from the remaining spectroscopically determined labels SPECTYPE=STAR and SPECTYPE=QSO (DESI Collaboration et al. 2024) along with the contaminants removed from the original SPECTYPE=GALAXY selections.
The predictions from each classifier for whether a source is a normal galaxy are simply the inverse of the boundary’s original predictions. As Fig. 12 shows, all classifiers capture nearly all galaxies across every magnitude bin. This behaviour is expected, as the selection criteria used by each classifier are explicitly designed to favour AGN-like features, thereby excluding typical galaxies by construction. The precision scores also match extremely closely between classifiers, indicating similar contaminants proving difficult, especially for the brighter sources. Such contaminants can be seen in Figs. 9 and 11, where stars and DESI-selected QSO fall outside the boundaries.
The ability to separate normal galaxies from other extragalactic or stellar objects benefits both groups of data. Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understanding of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al. 2025).
Stars
The star selection is made up of the positive instances of Euclid selected (star_candidate_euclid, MZ25), DESI-selected (SPECTYPE=STAR, DESI Collaboration et al. 2024), and Gaia-selected stars.
Due to the assumptions made for this method, the incidental capture of stars within our classifications is an expected outcome. The near-perfect recall shows that our classifier is able to capture stars well. This performance is beneficial when attempting to capture a high-purity galaxy sample, as discussed in the previous section. However, as other works such as Euclid Collaboration: Matamoro Zatarain et al. (2025) have shown, there are metrics to mitigate the contamination of stars. As a significant number of stars are likely to remain in our sample, for all subsequent selections, we apply the cut phz_star_prob_phz_class<0.3<0.3, to predictions of all classifiers to minimise the contamination of stellar sources.
Euclid-detected QSOs
@@
The 2-colour selection proved harder to match for all classifiers, with recall dropping to 0.6 for the faintest objects. Again, no significant differences are seen between the classifiers.
DESI-selected QSOs
@@
This selection provides the first evidence of significant differences in the suggested candidates for each classifier. The inpainting classifier is able to capture a significantly higher selection of lower magnitude sources compared to the other classifiers. Throughout all magnitude bands, the PSF and Srsic flux classifier has worse recall. The ROC classifier matched more candidates in the fainter sources.
Gaia-detected QSOs
For the Gaia-detected QSOs, we use the Quaia selection (Storey-Fisher et al. 2024). The subset consists of positive instances (qso_quaia, MZ25), with negative instances from Gaia-detected galaxies (in_galaxy_candidates_gaia, MZ25) and stars (star_candidate_gaia, MZ25). We again apply the MUMAX_MINUS_MAG $\leq-2.6$ restriction to the sources.
After the initial magnitude bin, our classifier matched with a high proportion of the Quaia-selected QSOs. The other two classifiers achieved equal recall scores, but in the faintest detected sources, our classifier was able to match with an additional 20% of sources.
@@
AllWISE R90 AGN
Utilising WISE-detected sources allows us to verify our ability to select candidates that may have been obscured by dust. Using the selection criteria derived by Assef et al. (2018), we compare our candidates to the 90% reliability (R90) selection, a high purity AGN sample.
@@
Much like the Gaia-detected QSOs, the small fraction of candidates compared to the full sample of possible sources leads to many of the classifiers selecting sources outside the R90 selection. With a completeness of only 17% (Assef et al. 2018) there will be many mid-infrared AGN in the negative sample, which the classifiers are likely finding.
@@
X-ray-selected AGN
We compare our candidates to the X-ray selection from the work by Euclid Collaboration: Roster et al. (2025). The positive instances are when the calculated probability that the source is a galaxy is sufficiently low (Gal_proba_roster <0.2<0.2, MZ25). The negative instances are sources with Gal_proba_roster $\geq 0.2$. Only sources with detection in X-rays are included in the results.
@@ 5.7 Decomposition of the AGN Component
Given our model’s ability to classify suspected AGN, it is not unreasonable for the reader to assume that the residuals of the repainting process and the output image can independently represent the AGN and galaxy components. Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals to accurately represent the AGN component for further scientific analysis has not been tested and, as such, cannot be used. A more significant exploration of which mask to use will also be vital for such a use case. A mask that adapts its size and shape for each input image will likely be necessary. For a deep-learning approach applying AGN decomposition on Euclid data, readers are referred to Euclid Collaboration: Margalef-Bentabol et al. (2025), hereafter called MB25, whose method provides AGN fraction ($F_{\text{AGN}}$) measurements that we use to analyse our AGN candidates.
The subset of the sources from the Q1 catalogue (Euclid Collaboration: Romelli et al. 2025) used within this work and the subset tested in MB25 feature an overlap of 425 781 sources. The confusion matrix for AGN predictions between the diffusion classifier and sources with an $F_{\text{AGN}}\geq 0.2$ (AGN candidate for MB25) are shown in Table 4. The majority of sources ($\sim 95\%$) are in agreement by both methods not to be AGN candidates. With the remaining sources, only $\sim 10\%$ agree, with over five times more AGN candidates selected by MB25. When comparing all the AGN candidates selected using their respective subsets of data, this number reduces down to 3.5 times more candidates selected by MB25 than the diffusion model (Table 5).
The left panel of Fig. 13 shows the AGN candidates selected by the diffusion classifier but predicted to have low AGN fraction by the model in MB25. The right panel shows the AGN fraction of sources that are agreed upon between the two methods. Fig. 26 shows where the MB25 selections are located on the inpainting metrics. As shown in Table 4, many of the MB25 sources are not located above the diffusion threshold. Of the sources that lie above the threshold, the majority show an $F_{\text{AGN}}>0.5$. This implies that the $F_{\text{AGN}}$ selection is more sensitive to fainter AGN.
Table 4: AGN candidate comparison between this work and the AGN fractions predicted by MB25. There is an overlap of 425 781 sources between the subset of Euclid Q1 sources selected by both methods.
@@
Table 5: Total number of AGN candidates from each morphology method and their respective data sets. The majority of AGN candidates from the two respective data sets are sources not used within the other’s data set.
@@
Figure 14: Comparison of our diffusion model selections with those from photometry-based selections (MZ25; Euclid Collaboration: Bisigello et al. 2024) and the selections from MB25. Although all the diffusion selected sources are plotted, to improve clarity, only a representative subset of the other selections is shown. The photometry-based selections are limited to much bluer sources, highlighted by the dotted line showing the cut introduced in MZ25.
Comparing the predictions for each of the Euclid-based selections in Fig. 14, we can see that our diffusion model covers areas of the colour space that the photometry-based selections do not. Being able to select both bluer galaxies, as captured by the photometry selections (MZ25; Euclid Collaboration: Bisigello et al. 2024), and the much redder sources captured by MB25, shows the versatility of the method, highlighting its effectiveness across different galaxy populations.
Figure 15: Impact of the diffusion noise scheduler on the median pixel values of the VIS images. The pixels covered in the median pixel calculation are shown as orange regions overlaid on the example source image in the leftmost plots. The top row shows how the additional noise impacts areas of the image that are more typically background noise, whereas the bottom row focuses on pixels that will be heavily influenced by the source. The binning of images according to their S/N shows how different levels of image clarity are affected by the gradual increase of introduced noise. Images with higher S/N can retain more of the source detail for more timesteps in the scheduler.
Figure 16: Top: Respective noised images produced by the cosine-beta schedule at different timesteps. Each image is a sample from the respective signal-to-noise bin directly below it. Due to the scales of pixel values, the introduced noise has a more significant impact on the typically fainter, low S/N images, leading to the images converging to Gaussian noise much sooner into the forward process. The relationship of S/N and rate of convergence results in the entirety of the top left of the grid of images being pure noise, indicating inefficient training for lower S/N images. This highlights the difficulty in applying off-the-shelf pipelines to the complexities of real-world astronomical data that feature high dynamic range and varying quality over images. Bottom: Distribution of S/N of galaxy images. Even though the sample is dominated by lower S/N images, a non-negligible number of sources with S/N¿1000 remains in the training set.
@@ 6.1 The effect of signal-to-noise
In machine-learning research, the improvements of generative models are often demonstrated with data sets that are of high resolution and typically noise-free. This provides a level of consistency across the full data set, ensuring the benchmarks reflect the model’s performance under ideal conditions. Whether through loss functions (as discussed in Sect. 2.3) or in the practical application of the pipelines, the model’s response to one part of the data will be similar to another.
However, the difficulty in learning the dynamics of our data is not consistent throughout the images. Higher signal-to-noise ratios and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low S/N counterparts. In the context of diffusion models, where the aim throughout training is to learn the dynamics of the introduced noise at different timesteps, the varying S/N of the images could again cause inconsistencies. This can be seen in both the classifier results shown in Fig. 12, where the fainter objects typically have lower recall, as well as the increased variance during inpainting of fainter images shown in Fig. 10.
To explore how the noise schedule impacts the Euclid images, Figure 15 shows the variability of pixel values as they progress through timesteps. We take a row and column of pixels (a slice) through each image, bin them by S/N, and compute the median pixel value at each index within the slice. This allows us to observe how the values change across the image. Since pixels affected by the source have much higher values than those near the edges of the image, Figure 15 compares slices through the outer regions (top) with slices through the centre of the image (bottom).
In the top row, we see that low S/N images are more significantly affected by the introduction of noise, and this effect appears earlier in the scheduler. The background noise converges to the Gaussian distribution across all S/N bins between halfway ($t=2000$) and three-quarters ($t=3000$) of the total noise steps. In contrast, for pixels dominated by the source (bottom row), high S/N images remain relatively unaffected through much of the schedule. The brightest images do not show noticeable changes until around $t=2000$, and only converge to the Gaussian distribution within the final 250 steps. It is evident that both the image’s S/N level and the position of the pixels within the image determine the number of timesteps needed for the noise to resemble the target Gaussian distribution.
@@
To improve the performance of a diffusion model, whether to better recreate a data set or to model more complex data, the simplest approach is to increase the number of unique denoising steps. The emphasis here is on uniqueness, as simply adding more timesteps can lead to redundancy, especially in the later stages where successive steps become too similar, add little value and increase overhead. This is precisely why Nichol & Dhariwal (2021) introduced the cosine-beta scheduler (discussed in Sect. 2.4); the original linear scheduler caused images to degrade into pure noise too early into the process.
The images in Fig. 16 show that for a large fraction of our data set, the model spends the majority of its time and computational effort learning the trivial task of denoising pure Gaussian noise, which provides no useful learning signal for galaxy morphology. This is a fundamental inefficiency in the implementation of diffusion-based models that not only affects Euclid images, but will apply to any scientific data set that exhibit varying levels of S/N. Although this has implications in regards to the efficiency and optimisation of the pipeline, it highlights the bigger issue of the data quality and variety assumptions implied when using off-the-shelf deep learning methods to the complex, heterogeneous data typical in large astronomical surveys and the wider scientific community. While a detailed implementation is beyond the scope of this work, this finding highlights a critical direction for future research: the development of adaptive noise schedulers that are conditioned on image properties like S/N. Such an advance would significantly improve both the efficiency and performance of diffusion models in scientific applications.
6.2 Considerations on training and inference costs
Our DDPM inpainting model required approximately 50 hours of training using a single Nvidia A6000 GPU, which aligns with expectations for models of similar scale and architecture. However, the inference pipeline is heavily constrained by memory rather than computation, as demonstrated by its near-linear scaling – the inpainting of 1024 and 8192 images took roughly the same amount of time when sufficient GPU memory was available. This memory constraint is due to the need to store and iteratively manipulate large intermediate tensors during inference. Training, although more computationally expensive, has a more structured process where gradients are computed and accumulated in a predictable manner.

### v3

@@ 5.6 Comparing to other classifiers
5.6.1 Srsic and PSF flux measurements
@@
To benchmark our non-parametric inpainting method against a more traditional approach, we constructed a classifier based on parametric modelling of galaxy light profiles. We utilised morphological parameters sourced directly from the Euclid MER Q1 catalogue (Euclid Collaboration: Romelli et al., 2025). Specifically, we used the provided flux_vis_sersic_euclid and flux_vis_psf_euclid measurements. These parameters were derived by the MER pipeline using the SourceExtractor++ software (Bertin et al., 2020; Kümmel et al., 2020). For each detected source, the pipeline performs two separate and independent model fits. The flux_vis_psf_euclid value represents the total flux obtained by fitting a model of the instrumental PSF to the source image. The flux_vis_sersic_euclid value represents the total flux from fitting a two-dimensional Srsic profile, which is designed to model the extended light distribution of a galaxy. The underlying assumption for a classifier based on these parameters is that sources with a dominant, unresolved central component (such as a QSO or a Type-1 AGN) will be described well by a PSF model, resulting in a high flux_vis_psf_euclid value. Conversely, normal resolved galaxies will be better described by the extended Srsic profile, leading to a high flux_vis_sersic_euclid value. By comparing the relative strength of these two flux measurements (see Fig. 15), we can construct a simple linear classifier to distinguish between point-like and extended sources. This baseline method is computationally inexpensive as it relies on pre-computed catalogue values, but it is fundamentally limited to this simple morphological distinction and may struggle with identifying AGN in complex host galaxies or at lower relative luminosities.
@@
The distribution of sources with these parameters is shown in Fig. 11. The equation of this linear boundary is $y=x^{1.13}$.
5.6.2 ROC metric
@@
Our second comparison is with the original ROC metric used in Sect. 4.3. As a clear distinction between the galaxy class and the others is visible, it is important to verify how such a metric performs as a classifier. This is especially true given that such a metric can be generated from the raw image, and no machine-learning model or precomputed values are required. The boundary for this classifier is $x=0.75$.
5.6.3 Performance against other candidates
@@
Given the variety and differences in the selection methods, there is an inherent incompleteness and potential for contamination. Therefore, our analysis focuses primarily on achieving sufficiently high recall. For an initial candidate selection method, such as the one proposed here, the primary goal is often to compile a large, inclusive sample for potential follow-up observations or more detailed multi-wavelength analysis. In this context, high recall is typically prioritised over high precision, which measures the purity of the selected sample. The scientific cost of missing a potentially interesting source (a false negative) is often considered higher than the operational cost of including a contaminant (a false positive) that can be filtered out in subsequent analysis stages. Therefore, our primary goal is to demonstrate the capability of our model to achieve high recall across a wide variety of AGN and QSO candidates which have been selected via different, more conventional methods and physical mechanisms.
Precision and recall are defined as follows:
@@
For a given classification task (e.g. identifying DESI QSO objects):
@@
The test set scores for each respective selection are shown in Fig. 12, with the collated results for stars, galaxies, AGN, and QSOs shown in Fig. 25. All predictions are made from images that were neither used during training nor used in the creation of any of the decision boundaries for the classifiers. Due to differing constraints for each selection, we split the analysis to allow for a more representative comparison.
5.6.4 Galaxies
Our galaxy selection was determined by the DESI spectral type classification SPECTYPE=GALAXY (DESI Collaboration et al., 2024), with AGN contaminants removed following the procedure detailed in Euclid Collaboration: Matamoro Zatarain et al. (2025), hereafter called MZ25. Negative instances were formed from the remaining spectroscopically determined labels SPECTYPE=STAR and SPECTYPE=QSO (DESI Collaboration et al., 2024) along with the contaminants removed from the original SPECTYPE=GALAXY selections.
The predictions from each classifier for whether a source is a normal galaxy are simply the inverse of the boundary’s original predictions. As Fig. 12 shows, all classifiers capture nearly all galaxies across every magnitude bin. This behaviour is expected, as the selection criteria used by each classifier are explicitly designed to favour AGN-like features, thereby excluding typical galaxies by construction. The precision scores also match extremely closely between classifiers, indicating that similar contaminants prove difficult, especially for the brighter sources. Such contaminants can be seen in Figs. 9 and 11, where stars and DESI-selected QSO fall outside the boundaries.
The ability to separate normal galaxies from other extragalactic or stellar objects benefits both groups of data. Being able to create a pure galaxy sample is vital for the precise measurements needed for cosmology, as well as to further our understanding of dark matter and dark energy, a primary object of the Euclid mission (Euclid Collaboration: Mellier et al., 2025).
5.6.5 Stars
The star selection is made up of the positive instances of Euclid selected (star_candidate_euclid, MZ25), DESI-selected (SPECTYPE=STAR, DESI Collaboration et al., 2024), and Gaia-selected stars. Due to the assumptions made for this method, the incidental capture of stars within our classifications is an expected outcome. The near-perfect recall shows that our classifier is able to capture stars well. This performance is beneficial when attempting to capture a high-purity galaxy sample, as was discussed in the previous section. However, as other works such as Euclid Collaboration: Matamoro Zatarain et al. (2025) have shown, there are metrics to mitigate the contamination of stars. As a significant number of stars are likely to remain in our sample, for all subsequent selections we applied the cut phz_star_prob_phz_class<0.3<0.3 to predictions of all classifiers to minimise the contamination of stellar sources.
5.6.6 Euclid-detected QSOs
@@
The two-colour selection proved harder to match for all classifiers, with recall dropping to 0.6 for the faintest objects. Again, no significant differences are seen between the classifiers.
5.6.7 DESI-selected QSOs
@@
This selection provides the first evidence of significant differences in the suggested candidates for each classifier. The inpainting classifier is able to capture a significantly higher selection of lower-magnitude sources compared to the other classifiers. Throughout all magnitude bands, the PSF and Srsic flux classifier has worse recall. The ROC classifier matched more candidates in the fainter sources.
5.6.8 Gaia-detected QSOs
For the Gaia-detected QSOs, we used the Quaia selection (Storey-Fisher et al., 2024). The subset consists of positive instances (qso_quaia, MZ25), with negative instances from Gaia-detected galaxies (in_galaxy_candidates_gaia, MZ25) and stars (star_candidate_gaia, MZ25). We again applied the MUMAX_MINUS_MAG $\leq-2.6$ restriction to the sources.
After the initial magnitude bin, our classifier matched with a high proportion of the Quaia-selected QSOs. The other two classifiers achieved equal recall scores, but in the faintest detected sources our classifier was able to match with an additional 20% of sources.
@@ 5.6.8 Gaia-detected QSOs
5.6.9 AllWISE R90 AGN
Utilising WISE-detected sources allows us to verify our ability to select candidates that may have been obscured by dust. Using the selection criteria derived by Assef et al. (2018), we compared our candidates to the 90% reliability (R90) selection, a high-purity AGN sample.
@@ 5.6.9 AllWISE R90 AGN
Much like the Gaia-detected QSOs, the small fraction of candidates compared to the full sample of possible sources leads to many of the classifiers selecting sources outside the R90 selection. With a completeness of only 17% (Assef et al., 2018) there will be many mid-infrared AGN in the negative sample, which the classifiers are likely finding.
@@
5.6.10 X-ray-selected AGN
We compared our candidates to the X-ray selection from the work by Euclid Collaboration: Roster et al. (2025). The positive instances are when the calculated probability that the source is a galaxy is sufficiently low (Gal_proba_roster <0.2<0.2, MZ25). The negative instances are sources with Gal_proba_roster $\geq 0.2$. Only sources with detection in X-rays are included in the results.
@@ 5.7 Decomposition of the AGN component
Given our model’s ability to classify suspected AGN, it is not unreasonable for the reader to assume that the residuals of the repainting process and the output image can independently represent the AGN and galaxy components. Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals in accurately representing the AGN component for further scientific analysis has not been tested and, as such, cannot be used. A more significant exploration of which mask to use will also be vital for such a use case. A mask that adapts its size and shape for each input image will likely be necessary. For a deep-learning approach applying AGN decomposition to Euclid data, readers are referred to Euclid Collaboration: Margalef-Bentabol et al. (2025), hereafter called MB25, whose method provides AGN fraction ($F_{\text{AGN}}$) measurements that we use to analyse our AGN candidates.
The subset of sources from the Q1 catalogue (Euclid Collaboration: Romelli et al., 2025) used within this work and the subset tested in MB25 feature an overlap of 425 781 sources. The confusion matrix for AGN predictions between the diffusion classifier and sources with an $F_{\text{AGN}}\geq 0.2$ (AGN candidate for MB25) are shown in Table 4. The majority of sources ($\sim 95\%$) are agreed by both methods not to be AGN candidates. For the remaining sources, only $\sim 10\%$ are agreed upon, with over five times more AGN candidates selected by MB25. When comparing all the AGN candidates selected using their respective subsets of data, this number reduces down to 3.5 times more candidates selected by MB25 than the diffusion model (Table 5).
The left panel of Fig. 13 shows the AGN candidates selected by the diffusion classifier but predicted to have a low AGN fraction by the model in MB25. The right panel shows the AGN fraction of sources that are agreed upon between the two methods. Fig. 26 shows where the MB25 selections are located on the inpainting metrics. As is shown in Table 4, many of the MB25 sources are not located above the diffusion threshold. Of the sources that lie above the threshold, the majority show an $F_{\text{AGN}}>0.5$. This implies that the $F_{\text{AGN}}$ selection is more sensitive to fainter AGN.
Table 4: AGN candidate comparison between this work and the AGN fractions predicted by MB25.
@@
Table 5: Total number of AGN candidates from each morphology method and their respective datasets.
@@
Figure 14: Comparison of our diffusion model selections with those from photometry-based selections (MZ25; Euclid Collaboration: Bisigello et al., 2024) and the selections from MB25. Although all the diffusion selected sources are plotted, to improve clarity, only a representative subset of the other selections is shown. The photometry-based selections are limited to much bluer sources, highlighted by the dotted line showing the cut introduced in MZ25.
Comparing the predictions for each of the Euclid-based selections in Fig. 14, we can see that our diffusion model covers areas of the colour space that the photometry-based selections do not. Being able to select both bluer galaxies captured by the photometry selections (MZ25; Euclid Collaboration: Bisigello et al., 2024) and the much redder sources captured by MB25 shows the versatility of the method, highlighting its effectiveness across different galaxy populations.
Figure 15: Impact of the diffusion noise scheduler on the median pixel values of the VIS images. The pixels covered in the median pixel calculation are shown as orange regions overlaid on the example source image in the leftmost plots. The top row shows how the additional noise impacts areas of the image that are more typically background noise, whereas the bottom row focuses on pixels that will be heavily influenced by the source. The binning of images according to their S/N shows how different levels of image clarity are affected by the gradual increase in introduced noise. Images with higher S/N can retain more of the source detail for more timesteps in the scheduler.
Figure 16: Top: Respective noised images produced by the cosine-beta schedule at different timesteps. Each image is a sample from the respective S/N bin directly below it. Due to the scales of pixel values, the introduced noise has a more significant impact on the typically fainter, low-S/N images, leading to the images converging to Gaussian noise much sooner into the forward process. The relationship between the S/N and the rate of convergence results in the entirety of the top left of the grid of images being pure noise, indicating inefficient training for lower-S/N images. This highlights the difficulty in applying off-the-shelf pipelines to the complexities of real-world astronomical data that feature a high dynamic range and varying quality over images. Bottom: Distribution of S/N of galaxy images. Even though the sample is dominated by lower-S/N images, a non-negligible number of sources with S/N¿1000 remains in the training set.
@@ 6.1 The effect of signal-to-noise
In machine-learning research, the improvements of generative models are often demonstrated with datasets that are of a high resolution and typically noise-free. This provides a level of consistency across the full dataset, ensuring the benchmarks reflect the model’s performance under ideal conditions. Whether through loss functions (as is discussed in Sect. 2.3) or in the practical application of the pipelines, the model’s response to one part of the data will be similar to another.
However, the difficulty in learning the dynamics of our data is not consistent throughout the images. Higher S/Ns and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low-S/N counterparts. In the context of diffusion models, where the aim throughout training is to learn the dynamics of the introduced noise at different timesteps, the varying S/N of the images could again cause inconsistencies. This can be seen in both the classifier results shown in Fig. 12, where the fainter objects typically have lower recall, as well as the increased variance during inpainting of fainter images shown in Fig. 10.
To explore how the noise schedule impacts the Euclid images, Fig. 15 shows the variability of pixel values as they progress through timesteps. We take a row and column of pixels (a slice) through each image, bin them by S/N, and compute the median pixel value at each index within the slice. This allows us to observe how the values change across the image. Since pixels affected by the source have much higher values than those near the edges of the image, Fig. 15 compares slices through the outer regions (top) with slices through the centre of the image (bottom).
In the top row, we see that low-S/N images are more significantly affected by the introduction of noise, and this effect appears earlier in the scheduler. The background noise converges to the Gaussian distribution across all S/N bins between halfway ($t=2000$) and three quarters ($t=3000$) of the total noise steps. In contrast, for pixels dominated by the source (bottom row), high-S/N images remain relatively unaffected through much of the schedule. The brightest images do not show noticeable changes until around $t=2000$, and only converge to the Gaussian distribution within the final 250 steps. It is evident that both the image’s S/N level and the position of the pixels within the image determine the number of timesteps needed for the noise to resemble the target Gaussian distribution.
@@
To improve the performance of a diffusion model, whether to better recreate a dataset or to model more complex data, the simplest approach is to increase the number of unique denoising steps. The emphasis here is on uniqueness, as simply adding more timesteps can lead to redundancy, especially in the later stages when successive steps become too similar, add little value, and increase overhead. This is precisely why Nichol & Dhariwal (2021) introduced the cosine-beta scheduler (discussed in Sect. 2.4); the original linear scheduler caused images to degrade into pure noise too early into the process.
The images in Fig. 16 show that for a large fraction of our dataset, the model spends the majority of its time and computational effort learning the trivial task of denoising pure Gaussian noise, which provides no useful learning signal for galaxy morphology. This is a fundamental inefficiency in the implementation of diffusion-based models that not only affects Euclid images, but will apply to any scientific dataset that exhibits varying levels of S/N. Although this has implications in regards to the efficiency and optimisation of the pipeline, it highlights the bigger issue of the data quality and variety assumptions implied when using off-the-shelf deep learning methods to the complex, heterogeneous data typical in large astronomical surveys and the wider scientific community. While a detailed implementation is beyond the scope of this work, this finding highlights a critical direction for future research: the development of adaptive noise schedulers that are conditioned on image properties such as S/N. Such an advance would significantly improve both the efficiency and performance of diffusion models in scientific applications.
6.2 Considerations for training and inference costs
Our DDPM inpainting model required approximately 50 hours of training using a single Nvidia A6000 GPU, which aligns with expectations for models of a similar scale and architecture. However, the inference pipeline is heavily constrained by memory rather than computation, as is demonstrated by its near-linear scaling – the inpainting of 1024 and 8192 images took roughly the same amount of time when sufficient GPU memory was available. This memory constraint is due to the need to store and iteratively manipulate large intermediate tensors during inference. Training, although more computationally expensive, has a more structured process whereby gradients are computed and accumulated in a predictable manner.